Polaski Company manufactures and sells a single product called a Ret. Operating at capacity, the company can produce and sell|-ဇာ
Answer #1

Requirement 1: Compute net increase in profits as follows

Particulars Price / Cost × Units = Amount
Sales Revenue ($47 × (1 − 16%)) (a) $39.48 × 8,000 = $315,840
Deduct: Costs
        Direct Material $20.00 × 8,000 = $160,000
        Direct Labor $6.00 × 8,000 = $48,000
        Variable Manufacturing Overhead $3.00 × 8,000 = $24,000
        Variable Selling Expense ($2 × (1 − 75%)) $0.50 × 8,000 = $4,000
        Special Machine ($16,000 ÷ 8,000 units) $2.00 × 8,000 = $16,000
                            Total Costs (b) $31.50 $252,000
Net Increase in Profits (a) − (b) $7.98 $63,840

Requirement 2: Compute net increase in profits as follows

Particulars Cost × Units = Amount
Costs Reimbursed
        Direct Material $20.00 × 8,000 = $160,000
        Direct Labor $6.00 × 8,000 = $48,000
        Variable Manufacturing Overhead $3.00 × 8,000 = $24,000
        Fixed Manufacturing Overhead $5.00 × 8,000 = $40,000
Fixed Fee $2.00 × 8,000 = $16,000
           Total Revenue (c) $36.00 $288,000
Deduct: Incremental Costs
        Direct Material $20.00 × 8,000 = $160,000
        Direct Labor $6.00 × 8,000 = $48,000
        Variable Manufacturing Overhead $3.00 × 8,000 = $24,000
                       Total incremental costs (d) $29.00 $232,000
Net Increase in Profits (c) − (d) $7.00 $56,000

Requirement 3: Compute net increase or decrease in profits as follows

Particulars Price / Cost × Units = Amount
Sales Revenue
              U.S Army $36.00 × 8,000 = $288,000
Deduct: Regular Channels $47.00 × 8,000 = $376,000
Decrease in Revenue ($88,000)
Deduct: Variable Selling Expenses avoided $2.00 × 8,000 = $16,000
Net decrease in profits if the order from Army accepted ($72,000)